2 blocks, M=5 kg &m=2kg are connected at the 2 ends of a spring of spring constant “k”=100N/m. Friction coefficient between ‘M’ & the ground is 0.4 & betwwen ‘m’ and the ground is 0.2.The minimum horizontal velocity that must be imparted to ‘m’ block towards right in order to just move ‘M’ block is?
(g=10m/s*s) 
 

Grade:12

1 Answers

Shobhit Varshney IIT Roorkee
9 years ago
Hi,

Suppose we impart velocity ‘v’ to the smaller block.
Now, applying law of conservation of energy,

0.5 mv2= f*x + 0.5 kx2

Now, kx = 0.4 * 50 * 10 (to just move the M)
So, x = 200 / 100 = 2m

putting this value of x in first equation,
0.5 * 2 * v2= 0.2 * 20 * 2 + 0.5 * 100 * 22

Calculate v

thanks.